WHAT IS THE FORMULAE TO FIND HEIGHT AND DISPLACEMENT OF A PARTICE AT TIME t????(in a projectile)

Grade:11

3 Answers

Aman Bansal
11 years ago

Dear Pritam

Formulae 

for height

 h = {v_0^2 \sin^2\alpha \over {2g}}

for displacement

 d = v_0 t_d \cos(\alpha) ,

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Rishab chhajer
11 years ago

Formulae to find height

USinΘt-0.5gt2

and formulae to find displacement is

UCosΘt

PRITAM DAS
11 years ago

Sir, 

the first formulae is independent of time then how can i get the height of a projectile at t secs....

I am not asking for maximum height.... rather the at a perticular point of time...